# Fundamental thresholds for computational and erasure errors via the coherent information

### v1

@@ Appendix B QEC codes
Appendix C The stabilizer configuration picture
@@ C.2 Calculation of $S(\rho_{RQ})$
C.3 Renyi coherent information
@@
where >delimited-⟨$\langle$..$\rangle$>⟨ … ⟩ > denotes erasure average. As expected, the disorder average over error chains is absent, however we can not get rid of the erasure average. Furthermore the erasure errors manifest themselves as infinite couplings that lock neighboring spins rigidly together. We can see this as fully depolarizing channels, i.e. $p=1/2$, producing $J\rightarrow\infty$ on the location of the erasures. Therefore one can just extend the existing mappings in the stabilizer configuration picture by adding fully depolarizing channels with some probability $e$. Besides, the contributions of erasure and computational errors are entangled in the average over different partition functions. In the absence of computational errors, $p=0$, the Hamiltonians have zero coupling yielding an infinite temperature state. Hence we get $2^{2k-h}$ inside the log and recover $I^{(r)}_{l}=(k-h_{l})\log 2$ for any $r$.

### v2

@@ Appendix B QEC codes
Appendix C Exact statistical mechanics model for [[15,3,3]] lift-connected surface code
In this section, we describe the exact statistical mechanics model for the lift-connected surface (LCS) code with parameters $L=3$ and $\ell=1$. To be concrete, we write down the exact form of the spin interactions; thus, the model applies to any choice of $p_{x}$, $p_{z}$, $p_{y}$, and $e$ (see Eq. (II.2)). Moreover, this serves as the starting point for constructing analogous models for arbitrary values of $L$ and $\ell$. First, let us note that there are two $\sigma$ and two $\tau$ spins per surface code sheet (see Fig. 12). Starting from Eq. (II.2), we now write down the terms $P_{\ell}^{X}$ for each qubit:
$P_{0}^{X}=\sigma_{L,0};\quad P_{1}^{X}=\sigma_{L,1};\quad P_{2}^{X}=\sigma_{L,2};\quad P_{3}^{X}=\sigma_{R,0};\quad P_{4}^{X}=\sigma_{R,1};\quad P_{5}^{X}=\sigma_{R,2};$ (77)
$P_{6}^{X}=\sigma_{L,0}\sigma_{L,2},\quad P_{7}^{X}=\sigma_{L,0}\sigma_{L,1};\quad P_{8}^{X}=\sigma_{L,1}\sigma_{L,2};\quad P_{9}^{X}=\sigma_{R,0}\sigma_{R,2};\quad P_{10}^{X}=\sigma_{R,0}\sigma_{R,1};\quad P_{11}^{X}=\sigma_{R,1}\sigma_{R,2};$ (78)
$P_{12}^{X}=\sigma_{L,0}\sigma_{R,0}\sigma_{R,1};\quad P_{13}^{X}=\sigma_{L,1}\sigma_{R,1}\sigma_{R,2};\quad P_{14}^{X}=\sigma_{L,2}\sigma_{R,2}\sigma_{R,0};$ (79)
We denote $\sigma_{L(R),q}$ as the spin on the left (right) belonging to the $q$-th surface code sheet. We identify three classes of interaction terms:
  • i)
    Single-spin terms [Eq. (77)] arising from the upper qubits on each surface code sheet.
  • ii)
    Two-spin terms [Eq. (78)] between spins on consecutive surface code sheets at the same in-sheet position, originating from the lower qubits on each surface code sheet.
  • iii)
    Three-spin terms [Eq. (79)] arising from the bulk qubits, i.e., those in the middle of the surface code sheet, where the two spins on the $q$-th sheet are coupled to the right spin on the $(q+1)$-th sheet.
At this point, it is worth noting that the spin interactions exhibit directionality along the vertical axis; that is, the interactions depend on vertical position, unlike the more symmetric interactions found in the Ising model derived from toric or surface codes.
Now we write down the $P_{\ell}^{Z}$ in terms of the $\tau$ spins in Fig. 12:
$P_{0}^{Z}=\tau_{U,0};\quad P_{1}^{Z}=\tau_{U,1};\quad P_{2}^{Z}=\tau_{U,2};\quad P_{6}^{Z}=\tau_{L,0};\quad P_{7}^{Z}=\tau_{L,1};\quad P_{8}^{Z}=\tau_{L,2};$ (80)
$P_{3}^{Z}=\tau_{U,0}\tau_{U,2},\quad P_{4}^{Z}=\tau_{U,0}\tau_{U,1};\quad P_{5}^{Z}=\tau_{U,1}\tau_{U,2};\quad P_{9}^{Z}=\tau_{L,3}\tau_{L,5};\quad P_{10}^{Z}=\tau_{L,3}\tau_{L,4};\quad P_{11}^{Z}=\tau_{L,4}\tau_{L,5};$ (81)
$P_{12}^{Z}=\tau_{U,0}\tau_{L,0}\tau_{L,1};\quad P_{13}^{Z}=\tau_{U,1}\tau_{L,1}\tau_{L,2};\quad P_{14}^{Z}=\tau_{U,2}\tau_{L,2}\tau_{L,0}.$ (82)
We define $\tau_{U(L),q}$ as the spin at the upper (lower) part of the $q$-th surface code sheet. The spin model obtained for the $Z$ stabilizers is simply a $90^{\circ}$ rotation of the spin model for the $X$ stabilizers. As a result, there is also directionality along the horizontal direction. To summarize, the three types of terms identified, namely the boundary terms [Eqs. (77), (78)] and the bulk terms [Eq. (79)]; serve as the building blocks for larger codes. The spin interactions exhibit a form of “chirality,” as they do not posses mirror symmetry along either spatial direction, even in the absence of disorder.
Appendix D The stabilizer configuration picture
@@ D.2 Calculation of $S(\rho_{RQ})$
D.3 Renyi coherent information
@@
where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes erasure average. As expected, the disorder average over error chains is absent, however we can not get rid of the erasure average. Furthermore the erasure errors manifest themselves as infinite couplings that lock neighboring spins rigidly together. We can see this as fully depolarizing channels, i.e. $p=1/2$, producing $J\rightarrow\infty$ on the location of the erasures. Therefore one can just extend the existing mappings in the stabilizer configuration picture by adding fully depolarizing channels with some probability $e$. Besides, the contributions of erasure and computational errors are entangled in the average over different partition functions. In the absence of computational errors, $p=0$, the Hamiltonians have zero coupling yielding an infinite temperature state. Hence we get $2^{2k-h}$ inside the log and recover $I^{(r)}_{l}=(k-h_{l})\log 2$ for any $r$.